OpenStudy (anonymous):

Calculate the restricted value, or values, for the equation below. (3/x+2) + (1/2x) =4/x+2

OpenStudy (dumbcow):

denominator can't be zero

OpenStudy (harsimran_hs4):

x cannot be equal to zero for the value of x you can solve this quadratic and by chance if you get 0 as answer you can leave it aside

OpenStudy (anonymous):

\[\frac{ 3 }{ x+2}+\frac{ 1 }{2x}=\frac{ 4 }{ x+2 }\] this is the equation

OpenStudy (dumbcow):

domain is restricted only function becomes undefined dividing by 0 is undefined set each denominator equal to zero x+2 = 0 ---> x = -2 2x = 0 ----> x=0 The restricted x-values are -2,0

OpenStudy (anonymous):

oh, okay. Thank you!